Together, two pipes can fill a swimming pool in 4 hours. Separately, the smaller pipe requires 6 more hours to fill the pool than the larger pipe. How many hours does it take for each pipe alone to fill the pool?
Answer by richwmiller(17219)

4/x+4/(x+6)=1
(8(x+3))/(x (x+6)) = 1
(8 (x+3))=(x (x+6))
8x+24=x^2+6x
x^2-2x-24
(x-6)*(x+4)=0
reject x=-4
x=6
one pipe takes 6 hours and the other takes 12
4/6+4/12=1
2/3+1/3=1